Please help, I'm stuck on this product and can't seem to figure out the answer.
directions: to simplify the following
4sqrt 9/8 - 4sqrt 2/9
the book says that the answer is 1/6 4sqrt 18. I have came close to this answer yet.
any help would be helpful please.
thanks, mike

Answer by stanbon(75887)

4sqrt 9/8 - 4sqrt 2/9
the book says that the answer is 1/6 4sqrt 18
-------------------------------
4*(3/2)sqrt(1/2) - (4/3)sqrt(2)
---
(6/2)sqrt(2) - (4/3)sqrt(2)
---
(3 - (4/3))sqrt(2)
--------------------
(5/3)sqrt(2)
==============================
Comment: This is not the answer your book has.
Why?
4sqrt 9/8 can be interpreted in several ways
I have interpreted as 4sqrt[(3^2/2^2*2) = 4(3/2)sqrt(1/2 = 6sqrt(2)/2 = 3sqrt(2)
But that does not result in the answer you have.
----
What to do?
Use parentheses so the reader can tell exactly what the problem is:
Is it 4sqrt(9)/8 or is it 4sqrt(9/8)?
Is it 2 sqrt(9)/2 or is ti 2 sqrt(9/2)?
--------------------
Cheers,
Stan H.